OpenStudy (anonymous):

NUMBER THEORY: Show that 2938572369 is or is not divisible by 11. @jim_thompson5910 @hartnn @TuringTest @campbell_st @robtobey @RadEn @Luigi0210 @wio @dan815 @Compassionate @Loser66 @shamil98 @Preetha @Opcode @april115 @jagatuba @AllTehMaffs @wolfe8 @SnuggieLad @tyteen4a03 @gypsy1274 @petiteme @Gatorgirl @AngelWilliams16 @DSS @turtleluver @lopezking1 @jojo4eva @Nuke_ur_friend @Kristen17 @Smexi_Girl @shelbygt520 @sunnyshores @Idealist @caitlinnr14 @forevershorty @lovatic4life @AccidentalAiChan @Victoria_Davis_2018

OpenStudy (anonymous):

Please do not tag a lot of people at one time

OpenStudy (shamil98):

Just use long division.

OpenStudy (anonymous):

I have to use number theory principles to evaluate. I was never taught the 11s trick though.

OpenStudy (anonymous):

Yolo.

OpenStudy (jagatuba):

First add all the even placed numbers. Then add all the odd placed numbers. Subtract the even sum from the odd sum. If the result is 0 or a multiple of 11 then the original number is divisible by 11.

OpenStudy (shamil98):

don't know the number theory principles, but through long division this is what i got :S

OpenStudy (shamil98):

was going to just use the draw tool on here, but not enough space..

OpenStudy (raden):

2938572369 the sum of odd position numbers : (2+3+5+2+6) = 18 the sum of even position numbers : (9+8+7+3+9) = 36 36-18=18 18 divisible 11 or not ?

OpenStudy (anonymous):

That. Was awesome.

OpenStudy (jagatuba):

Example: 9867 (8+7)-(9+6) 15-15=0 9867 is divisible by 11

OpenStudy (anonymous):

\[2938572369=3^2*3583*91127 \]

OpenStudy (anonymous):

\[\text{Mod}[2938572369,11]=7 \]

OpenStudy (jagatuba):

Yeah @robtobey. I'm not sure the relevance to the question, but as long as you went off topic, we can also show: